At the PS, replacement of the old cooling and ventilation system, which dated back to 1957, is progressing well. The dismantling phase has just been completed and the tunnel is being prepared to receive a brand new ventilation system offering more efficient aeration of radioactive areas. At the same time, testing of the newly installed access system is under way.
At the SPS, all the equipment of the Long Straight Section 1 (LSS1) has been removed to make way for the campaign to replace the irradiated cables, which started this week.
In the LHC, the Radiation to Electronics (R2E) campaign is still progressing at a good clip, especially at Point 1, where the teams are two and a half months ahead of schedule.
The Superconducting Magnets and Circuits Consolidation (SMACC) project is going entirely to plan, with the outer sleeves (W) having been removed from the equivalent of seven sectors and leak tests already in progress in several sub-sectors. Two sub-sectors have passed and another five are on the point of being tested. The installation of the shunts has also been progressing well, with one third already in place.
Two of the DFBA electrical feedboxes have already been consolidated, but three faulty ones have been identified, with damaged gimbal bellows that will need to be replaced. One of these replacements is under way and the second will commence shortly. However, the third one, at LHC Point 6, is proving difficult to access and part of it will probably have to be brought up and worked on at the surface.
Nearly all the 1,344 DN200 safety valves designed to release the helium in the event of pressure build-ups in the accelerator have now been installed. This work should be completed by the end of October.
The X-ray testing campaign in the tunnel, aimed at detecting faults in the machine’s cryogenic distribution system, is ahead of schedule with 60% of the tests already completed.
Finally, 12 out of the 18 magnets replaced in the accelerator have already been reconnected.